A confined quantity of gas is at a pressure of 2.50 atm and a temperature of -10°C, what is the pressure if the temperature is increased to 22°C

1
Expert's answer
2022-04-16T04:09:27-0400

The process described is called an isochoric process or constant-volume process. It can be characterized by the following equation:

"P\/T = const"

That means:

"P_1\/T_1 = P_2\/T_2" , where 1,2 -- states before and after temperature increase and temperature is expressed in K.

Find P2:

"P_2 = P_1 * T_2 \/ T_1 = 2.5 * 295\/263 = 2.80 (atm)"

Answer: 2.80 atm is the pressure of the gas after its temperature was increased.